OpenStudy (anonymous):

Expand into equivalent log expression: In x(x+1)

OpenStudy (anonymous):

ln(x)+ln(x+1)

jimthompson5910 (jim_thompson5910):

jahtoday is correct. He's using the identity \[\ln(A+B)=\ln(A)+\ln(B)\]